Find the value of (g • f)(7) if f(x)=3x-4 and g(x)=3x+1...
Answer by nerdybill(7384)

Find the value of (g • f)(7) if f(x)=3x-4 and g(x)=3x+1
.
(g • f)(x) = (3x-4)(3x+1)
FOIL the right side:
(g • f)(x) = 9x^2+3x-12x-4
(g • f)(x) = 9x^2-9x-4
.
finally:
(g • f)(7) = 9(7)^2-9(7)-4
(g • f)(7) = 9(49)-63-4
(g • f)(7) = 441-63-4
(g • f)(7) = 378-4
(g • f)(7) = 374